OpenStudy (anonymous):

Find the exact area of a circle having the given circumference. 4(sqrt3 PI )

OpenStudy (amistre64):

Do you recall the formula for the area of a circle?

OpenStudy (anonymous):

um no i forgot sorry

OpenStudy (amistre64):

can you look it up for me? its kind of needed for an answer i believe ..

OpenStudy (anonymous):

i think its A =sqrt r2

OpenStudy (amistre64):

so close, lets try A = pi r^2 tho, that might work better ;)

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

now if we knew "r" we could find the answer; what do you remember about the formula for the circumference of a circle?

OpenStudy (anonymous):

Amistre, you're doing a good deed helping kids out on this site. Just wanted to let you know I appreciate it, even though this question doesn't pertain to me.

OpenStudy (anonymous):

im not sure im home schooled and im rlly tryna learn all by myself

OpenStudy (amistre64):

thnx hood. Johny, your doing fine, the formula for the circumference of a circle will give us the "r" we need for the area. But i need you to recall something about that formula, anything at all would be helpful, right or wrong.

OpenStudy (anonymous):

THEOREM 7-8 The area of a circle is the product sqrt of and the square of the radius of the circle: A =sqrt r2. Given: Circle P with radius r and circumference C. this is all i was given

OpenStudy (amistre64):

The circumference of a circle has the formula, and this is what sucks about math at this level is simply having to remember the formulas, is: C = d pi , where d = diameter it can also be stated as C = pi 2r, or C = 2pi r

OpenStudy (amistre64):

they give you: 4sqrt3 PI as the circumference so lets plug that in for C and see if we can find "r" 4sqrt3 pi = 2pi r ; divide off the pi from both sides 4sqrt = 2r ; now divide off the 2 from both sides 2sqrt = r and this is our "r" value

OpenStudy (anonymous):

ok so is that my answer

OpenStudy (amistre64):

this is how we get to the answer ... we are almost there

OpenStudy (amistre64):

Do you understand how I found "r" from the information they gave us?

OpenStudy (amistre64):

i seem to have typoed out my "3" r = 2sqrt(3) is better

OpenStudy (anonymous):

yes i understand

OpenStudy (amistre64):

good, becasue the process is the same regardless of the numbers they give us. So as long as we understand the process we can solve any numbers thrown our way.

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

now, that we have an "r" value, lets find the Area: A = pi r^2 A = pi (2sqrt(3))^2 A = pi * 2^2 * sqrt(3)^2 A = pi *4 *3 A = pi *12 and to clean it all up Area = 12pi

OpenStudy (amistre64):

as long as i did the math right .... which has its good days and its bad ones .... that whould do it

OpenStudy (anonymous):

oooo i seeeee

OpenStudy (amistre64):

it takes practice, and remembering formulas and such, but its all good in the end :) good luck with your self teachings; im self taught as well

OpenStudy (anonymous):

lol ok thnx